Determine the Green, s function and express the solution as a definite integral

-y''=f(x), y(0)=0, y'(1)=0


1
Expert's answer
2021-09-10T09:24:09-0400

Answer:-

Given IVP:



"-y''(x)=f(x),\quad y(0)=0,\quad y(1)=0"

The general solution of homogeneous equation


"y(x)=c_1+c_2x"

Hence,


"G(x,s)=\left\{\begin{matrix}
 c_1+c_2x,\quad 0\leq x\leq s, \\
 c_3+c_4x,\ \quad s< x\leq 1.
\end{matrix}\right."

Initial conditions give


"G(0,s)=c_1=0,\quad G(1,s)=c_4=-c_3"

So


"G(x,s)=\left\{\begin{matrix}
 c_2x,\quad 0\leq x\leq s, \\
 c_3(1-x),\ \quad s< x\leq 1.
\end{matrix}\right."

The continuity of Green function gives


"G(x,s)=\left\{\begin{matrix}
 x(s-1),\quad 0\leq x\leq s, \\
 s(x-1),\ \quad s< x\leq 1.
\end{matrix}\right."

The solution of DE


"y(x)=\int_0^1G(x,s)f(s)ds""=\int_0^xx(s-1)f(s)ds+\int_x^1s(x-1)f(s)ds"